OpenStudy (anonymous):

I need a little help with finding the derivative. if you could please explain step by step. I would really appreciate it. Find the derivative of f(x) = 5/x at x = -1

OpenStudy (anonymous):

i am only in pre-calc btw

OpenStudy (anonymous):

just in case you use a calc shortcut

OpenStudy (jdoe0001):

not sure ... how to cover it then.... because using Limits, would be calc, and using the power rule will also be calc.....

OpenStudy (jdoe0001):

hmmm well.. actually limits would be precalc so

OpenStudy (anonymous):

what was included in the lesson was (f(x+h)-f(x))/h

OpenStudy (anonymous):

to find derivative

OpenStudy (jdoe0001):

\(\bf lim_{x\to -1}\quad \cfrac{5}{x}\) will do IMO

OpenStudy (anonymous):

there are answer choices though

OpenStudy (anonymous):

and thats not the derivative

OpenStudy (anonymous):

im just not sure how to move everything out of the numerator's denominator

OpenStudy (anonymous):

have you been taught how to do \[f'(a) = \lim_{x \rightarrow a} \frac{ f(x)- f(a) }{ x-a }\]

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

\(\bf lim_{x\to -1}\quad \cfrac{5}{x}\implies lim_{h\to 0}\quad \cfrac{f(x+h)-f(x)}{h}\\ \quad \\ \implies lim_{h\to 0}\quad \cfrac{\frac{5}{x+h}-\frac{5}{x}}{h}\implies lim_{h\to 0}\quad \cfrac{\frac{5x-5x-5h}{x(x+h)}}{h}\\ \quad \\ lim_{h\to 0}\quad \cfrac{\frac{-5h}{x(x+h)}}{h}\implies lim_{h\to 0}\quad \cfrac{-5h}{x(x+h)}\cdot \cfrac{1}{h}\implies \cfrac{-5}{x(x+h)}\)

OpenStudy (jdoe0001):

and as you can see, once h = 0, then we end up with \(\bf \implies \cfrac{-5}{x(x+0)}\implies \cfrac{-5}{x^2}\)

OpenStudy (anonymous):

i thought that you cant set h to 0 when its in the denominator

OpenStudy (jdoe0001):

well, yes, you're correct IF it makes the rational undefined, after the factoring and simplifying it does not